Originally, the dimensions of a rectangle were 18cm by 24cm. When both dimensions were decreased by the same amount, the area of the rectangle decreased by 117 sq. cm. Find the dimensions of the new rectangle.
Answer by mananth(16949)

rectangle dimensions
18 by 24 cm
Area = 18*24=432 cm^2
..
let them be decreased by x
..
new dimensions =
18-x
24-x
..
area = (18-x)(24-x)
area = 432-18x-24x+x^2
area=432-42x+x^2
..
Original area - reduced rectangle area = 117
432-(432-42x+x^2)=117
432-432+42x-x^2=117
42x-x^2=117
x^2-42x+117=0
x^2-39x-3x+117=0
x(x-39)-3(x-39)=0
(x-3)(x-39)=0
So x= 3. x= 39 is not practical.
dimensions of new rectangle = 15cm by 21cm
..
m.ananth@hotmail.ca